Suppose that the pressure in an automobile tire is 2.30 atm at a tenperature of 20.0 degree celsius. What will the pressure in the tire be if after 10 miles of driving the temperature of the tire increases to 47.0 degree celsius
1
Expert's answer
2019-08-20T04:55:31-0400

T1 = 293.15 K; T2 = 320.15 K.

V = const.

P1/T1 = P2/T2

2.30 atm / 293.15 K = P2 / 320.15 K

P2 = 2.30 atm × 320.15 K / 293.15 K = 2.5118 atm ≈ 2.51 atm.